In two hours a boat is rowed downstream 4 miles and back again. If the speed of the stream is 1.5 kts what is that of the boat?
Found 2 solutions by scott8148, Alan3354:
Answer by scott8148(6628)

[4 / (b + 1.5)] + [4 / (b - 1.5)] = 2

4b - 6 + 4b + 6 = 2 (b^2 - 2.25)

0 = b^2 - 4b - 2.25

use quadratic formula to find b


miles and kts are NOT in the same units
___ this solution does not contain a correction for the difference

Answer by Alan3354(69443)

miles and kts are NOT in the same units
-----------
miles is distance, knots is speed.